The perimeter of a rectangular field is 606m.The length is 42 more than twice the width. Find the dimension.

Answer by ewatrrr(24785)


Hi,
Let x and (2x +42cm)represent the width and length respectively

Question States***
2(2x+42cm) + 2x = 606m
Solving for x
4x + 84cm +2x = 606cm
6x = 522cm
x = 87cm, the width. Length is 216cm. (2*87cm +42cm)
CHECKING our Answer***
2*87cm + 2*216cm = 606m